OpenStudy (anonymous):

Exponents and logarithms and radical numbers...sooo simplify the expression nr.2

OpenStudy (anonymous):

\[2\sqrt{8}+\sqrt{200}-4\sqrt{18}\]

OpenStudy (anonymous):

The only way you can simplify is after you simplify the radicals.

OpenStudy (anonymous):

\[ 2\sqrt{8}=2\sqrt{2\cdot 2\cdot 2}=2\cdot 2\sqrt{2}=4\sqrt{2} \]

OpenStudy (anonymous):

\[ \sqrt{200} = \sqrt{2\cdot 2\cdot 2\cdot 5\cdot 5}=5\cdot 2\sqrt{2}=10\sqrt{2} \]

OpenStudy (anonymous):

Now you need to do the same with \[ 4\sqrt{18} \]But do it yourself this time for practice.

OpenStudy (anonymous):

yeah I got the same thing,but I don't know how to do with \[4\sqrt{18}\] should it be \[4\sqrt{9}\times \sqrt{2}\]? and soo? @wio

OpenStudy (anonymous):

Okay what is the prime factorization of \(18\)?

OpenStudy (anonymous):

Oh wait, what you have is fine, but you need to change \(\sqrt{9}=\sqrt{3\cdot 3}=3\)

OpenStudy (anonymous):

ohh thanks that's why I got the wrong answer hehe ^^